Shelly deposits $9200 at 5% interest semiannually find the amount after 2 pay periods of interest?
Found 2 solutions by swillauer04, MathTherapy:
Answer by swillauer04(1)

Set up a proportion x/9200= 5/100, simplify 5/100 to 1/20. Then, cross multiply the proportion x/9200= 1/20. You will get the equation 20x= 9200. Next, divide 20x by 20 to cancel out the coefficient, and using the division property of equality, divide 9200 by 20 as well. You will get the equation x= 460. Because you are trying to find a new value after depositing an initial value of 9200 and 2 interest payments of 460, you will get a new value/ sum of 10,120.

Answer by MathTherapy(10551)


Shelly deposits $9200 at 5% interest semiannually find the amount after 2 pay periods of interest?
With semi-annual compounding interest: $9,665.75
With semi-annual simple interest: $9,660.00